A ball (m=2.50kg) rolling across the ground experiences a decrease in its kinetic energy from 200.0 J to 60.0 J. Calculate the work lost due to friction. What was the initial and final velocity?
1
Expert's answer
2018-11-27T09:28:10-0500

The work lost

W=KE_f-KE_i=60.0-200=-140 J

Since

KE=(mv^2)/2

we get

v_i=√((2KE_i)/m)=√((2×200)/2.50)=12.6 m/s

v_f=√((2KE_f)/m)=√((2×60.0)/2.50)=6.93 m/s